OpenStudy (anonymous):

how long will a 25 kg block to fall to the ground 78.4 m below

OpenStudy (anonymous):

v=d/t

OpenStudy (anonymous):

f=ma so 25kg (9.8m/s/s)=245N then use the another formula that works force and distance and velocity together.

OpenStudy (anonymous):

d=1/2 * a * t2 (t square) so 78.4=1/2*g*t2 78.4=1/2*9.8*t2 t2=16 t=4 sec